# Edge of Many-Body Quantum Chaos in Quantum Reservoir Computing

### v1

@@ Abstract
Reservoir computing (RC) is a machine learning paradigm that harnesses dynamical systems as computational resources. In its quantum extension—quantum reservoir computing (QRC)—these principles are applied to quantum systems, whose rich dynamics broadens the landscape of information processing. In classical RC, optimal performance is typically achieved at the “edge of chaos,” the boundary between order and chaos. Here, we identify its quantum many-body counterpart using the QRC implemented on the celebrated Sachdev-Ye-Kitaev model. Our analysis reveals substantial performance enhancements near two distinct characteristic “edges”: a temporal boundary defined by the Thouless time, beyond which system dynamics is described by random matrix theory, and a parametric boundary governing the transition from integrable to chaotic regimes. These findings establish the “edge of many-body quantum chaos” as a fundamental design principle for QRC.
@@
More recently, the power of quantum has also integrated with reservoir computing (RC), a machine learning framework renowned for its efficacy in time-series analysis [12, 13, 14, 15, 16]. True to its name, the most distinctive feature of RC is the use of an input-driven dynamical system, known as a “reservoir”, to nonlinearly map input data onto a feature space. Notably, training in RC is limited to the post-processing stage, and the reservoir itself remains fixed. Extending this idea into the quantum domain, Fujii and Nakajima introduced quantum reservoir computing (QRC) [17]. QRC harnesses natural quantum dynamics as a reservoir, generating rich, nonlinear mappings in an exponentially large Hilbert space without encountering the trainability bottlenecks prevalent in VQAs. This capability has sparked significant theoretical progress [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], as well as promising experimental demonstrations [37, 38, 39, 40, 41, 42, 43, 44].
@@
Given the optimization-free nature of RC, researchers naturally ask: What characterizes an effective reservoir system? While this question remains largely unanswered for QRC, several key criteria have been identified for classical RC. The most renowned among them is the concept of the “edge of chaos” [45, 46, 47, 48, 49]. This principle posits that the boundary between chaotic and nonchaotic dynamical regimes provides a performance sweet spot for RC systems, as empirically verified across diverse platforms [50, 51, 52]. By analogy, exploring a quantum version of the edge of chaos offers a promising route to addressing the fundamental question posed above.
@@
In classical systems, chaos is naturally framed via phase space trajectories [53]. In quantum mechanics, however, the Heisenberg uncertainty principle precludes such a direct picture, and one must instead rely on more subtle diagnostics. Early investigations of quantum chaos therefore focused on single-particle systems whose classical limits are chaotic, leading to the conjecture (with some analytical supports) that their spectral statistics are governed by random matrix theory (RMT) [54, 55, 56, 57, 58]. In recent years, attention has shifted to many-body settings, spurred in large part by the exactly solvable Sachdev-Ye-Kitaev (SYK) model [59, 60, 61, 62]. Various diagnostics, including level statistics [63, 64, 65, 66], out-of-time-order correlators [67, 68, 69], and spectral form factors [70, 71, 72, 73] have enriched our understanding of chaos in interacting quantum systems. Experimental realizations of the SYK physics have also been proposed on several platforms [74, 75, 76, 77, 78]. Notably, the SYK model exhibits maximally quantum chaotic behavior, saturating the so-called bound on chaos [79, 69, 80], and has thus stood as a canonical model for the investigation of many-body quantum chaos. These insights motivate the exploration of QRC built upon the SYK model, aiming to address the open question of reservoir effectiveness from a quantum chaotic standpoint.
@@
In this Letter, we explicitly demonstrate key characteristics of effective quantum reservoir systems. In particular, we examine the QRC performance near two distinct chaotic boundaries of the SYK model. The first is a temporal boundary defined by the Thouless time, beyond which the system’s dynamics conforms to RMT. The second is a parametric boundary marking the transition from quantum chaotic to integrable (nonchaotic) behavior, controlled by the strength of an additional noninteracting term. Using representative machine learning tasks that demand both memory retention and nonlinear transformation, we demonstrate that QRC performance is maximized near the onset of quantum chaotic regime in both temporal and parametric domains. These observations establish the “edge of many-body quantum chaos” as a foundational design principle for QRC systems.
@@
where $N$ denotes the system size, and $c^{\dagger}_{i}$ ($c_{i}$) is the spinless fermionic creation (annihilation) operator at site $i$. These operators satisfy the anticommutation relations $\{c_{i},c_{j}^{\dagger}\}=\delta_{i,j}$ and $\{c_{i},c_{j}\}=\{c_{i}^{\dagger},c_{j}^{\dagger}\}=0$. In Eq. (1), the first term describes the SYK4 interaction: the all-to-all couplings $J_{ijkl}$ are drawn from a zero-mean complex Gaussian distribution of variance $\overline{|J_{ijkl}|^{2}}=J_{4}^{2}/N^{3}$ ($\overline{\bullet}$ denotes the ensemble average). The second term is the noninteracting SYK2 contribution, with coupling constants $\kappa_{ij}$ sampled from a zero-mean complex Gaussian distribution of variance $\overline{|\kappa_{ij}|^{2}}=\kappa_{2}^{2}/(2N)$. The couplings satisfy the symmetry relations $\kappa_{ij}=\kappa_{ji}^{*}$, $J_{ijkl}=J_{klij}^{*}$, and $J_{ijkl}=-J_{jikl}=-J_{ijlk}$. Both coupling terms conserve the total fermion number $N_{p}\equiv\sum_{i}\langle c_{i}^{\dagger}c_{i}\rangle$. Unless otherwise specified, we set $N=8$, $J_{4}=1$, and $N_{p}=N/2$ (note that the QRC input protocol introduced below does not conserve $N_{p}$).
@@
To diagnose the chaotic nature of quantum many-body systems, we first examine the level statistics [63, 64, 65, 66]. Let the ordered eigenvalues of $\mathcal{H}$ be $E_{n}$; the spacings between consecutive energy levels are given by $s_{n}=E_{n+1}-E_{n}$. In an integrable system, the spacing distribution follows Poisson statistics, whereas in a quantum chaotic system, it is well described by Wigner-Dyson (WD) distribution. A convenient metric to characterize these distributions is the level spacing ratio, defined as $r_{n}=\min(s_{n}/s_{n+1},s_{n+1}/s_{n})$. Its mean value $\langle r\rangle$ assumes $\approx 0.3862$ for Poisson distribution and $\approx 0.5996$ for WD statistics [63, 64]. Figure 2(a) presents the distribution of the level spacing ratio $P(r)$ for the SYK4 model [$(J_{4},\kappa_{2})=(1,0)$] and the SYK2 model [$(J_{4},\kappa_{2})=(0,1)$], in excellent agreement with the WD and Poisson predictions, respectively. Notably, when both terms are present, the system can be tuned between these extremes: as the ratio of the typical coupling strengths $\kappa_{2}/J_{4}$ increases, Fig. 2(b) displays a gradual shift of $\langle r\rangle$ from the WD to the Poisson value. Despite the smearing effects associated with finite system sizes, this result strongly suggests the emergence of the quantum chaotic to integrable transition [81, 82, 83, 84, 85] (see the Supplemental Material [86]).
@@
where $\mathrm{Tr}_{1}\{\bullet\}$ denotes the partial trace over site $1$; see the Supplemental Material [86] for details on the observed dynamics. During the evolution, the expectation value of the single-site particle number is measured at discrete times $t_{v}=v\Delta t_{\mathrm{in}}/V$, where $0<v\leq V$ and $V$ is the number of virtual computational nodes (here, $V=10$). At the $k$-th step, the outcomes $\mathrm{Tr}\{\rho^{(k)}(t_{v})c_{i}^{\dagger}c_{i}\}$ are assembled into a one-dimensional state vector $\bm{\mathrm{x}}^{(k)}$; including an additional constant term, $\bm{\mathrm{x}}^{(k)}$ has $NV+1$ components in total. The output layer then applies a linear transformation $y^{(k)}=W_{\mathrm{out}}\bm{\mathrm{x}}^{(k)}$, where the output weight $W_{\mathrm{out}}$ is trained by minimizing the mean squared error between the computed output $y^{(k)}$ and the target output $\bar{y}^{(k)}$. In our simulations, a total of $10000$ inputs are processed. Starting from a random initial state, the first $4000$ inputs are discarded to wash out the effects of initial conditions; for the convergence property, see the Supplemental Material [86]. The subsequent $3000$ inputs are employed for training $W_{\mathrm{out}}$, and the final $3000$ are reserved for evaluation. Performance results are averaged over $500$ independent random realizations.
We evaluate the reservoir performance on two complementary tasks: the short-term memory (STM) task, which probes linear memory capacity, and the nonlinear auto-regressive moving average (NARMA) task, which evaluates combined memory and nonlinear processing capabilities [14, 17].
@@
In the STM task, the target output at delay $d$ is defined as the input value $d$ steps earlier in a random input sequence $\{u^{(k)}\}$, i.e., $\bar{y}^{(k)}_{d}=u^{(k-d)}$. During the testing phase, we collect the outputs $\bm{\mathrm{y}}=\{y^{(k)}\}$ and targets $\bar{\bm{\mathrm{y}}}_{d}=\{\bar{y}^{(k)}_{d}\}$, and compute the determination coefficient $R^{2}_{d}\equiv\mathrm{cov}^{2}(\bm{\mathrm{y}},\bar{\bm{\mathrm{y}}}_{d})/[\sigma^{2}(\bm{\mathrm{y}})\sigma^{2}(\bar{\bm{\mathrm{y}}}_{d})]$, where $\mathrm{cov}$ and $\sigma^{2}$ denote the covariance and the variance, respectively. $R^{2}_{d}$ is bounded between 00 and $1$; values closer to $1$ indicate better memory performance.
[…]
@@
Temporal edge of many-body quantum chaos. —Our investigation begins with an examination of the QRC performance of the SYK4 and SYK2 models, in which clear fingerprints of many-body quantum chaos manifest themselves in the temporal domain. As a time-domain probe of the energy-spectrum correlations characteristic of quantum chaos, we analyze the spectral form factor (SFF) $K(t)=\overline{\sum_{m,n}e^{i(E_{m}-E_{n})t}}/\mathcal{N}^{2}$, where $\mathcal{N}$ denotes the Hilbert space dimension [70, 71, 72, 73]. In many-body chaotic systems, the SFF generally exhibits a “ramp-plateau” profile governed by two time-scales: the Thouless time $t_{\mathrm{Th}}$, after which the evolution shows a universal “ramp” and follows RMT predictions, and the plateau time $t_{\mathrm{p}}$, beyond which $K(t)$ saturates at a plateau. Figures 3(a) and 3(b) illustrate the SFF for the SYK4 and SYK2 limits, respectively. Notably, in the SYK4 case, the SFF exhibits the pronounced ramp beginning around $t_{\mathrm{Th}}\sim 4.5$ and the subsequent plateau beyond $t_{\mathrm{p}}\sim 80$, providing a clear signature of many-body quantum chaos. While $K(t)$ in the SYK2 model also features a dip and a plateau, they occur on a much shorter timescale and with a different system-size scaling of $t_{\mathrm{p}}$ compared to the chaotic system (see the Supplemental Material for details [86]).
@@
Figure 3: (a), (b) The SFF $K(t)$ for the SYK4 and SYK2 models, with each curve averaged over $20000$ realizations. (c), (d) The QRC performances as functions of $\Delta t_{\mathrm{in}}$ for the SYK4 and SYK2 models. The upper panels show the memory performance $R^{2}_{d=n-1}$ (higher values indicate better retention), while the lower panels display the NMSE on the NARMA tasks (lower values indicate better accuracy). Markers represent the orders $n=2,3,5$, and $7$ (plotted sparsely for brevity). The horizontal dotted lines indicate the reference performance of the Haar QRC model, with poorer performance corresponding to higher $n$. Performance results are averaged over $500$ realizations, with shaded regions denoting the standard deviation. Vertical dashed lines in (a) and (c) represent the Thouless time $t_{\mathrm{Th}}$.
@@
Turning to computational benchmarks, Figs. 3(c) and 3(d) plot the QRC performance of the SYK4 and SYK2 models as functions of the input interval $\Delta t_{\mathrm{in}}$, which defines the intrinsic time-scale governing the QRC dynamics. The upper panels display the STM performance $R^{2}_{d=n-1}$ at delay $d=n-1$, and the lower panels show the NMSE on the NARMA-$n$ tasks of orders $n=2$, $3$, $5$, and $7$. The chosen STM delay corresponds to the oldest input term explicitly appearing in the recursive definition of the NARMA-$n$ dynamics; additional past inputs are nevertheless implicitly encoded within the target sequence $\{\bar{y}^{k}\}$ (see the Supplemental Material for the full delay dependence of $R^{2}_{d}$ [86]). To establish a baseline for comparison, we also analyze the Haar QRC model, in which the evolution operator $e^{-i\mathcal{H}t}$ is replaced by a unitary $U$ drawn from the Haar measure [87, 88]. Its performance, shown as horizontal dashed lines, deteriorates monotonically with increasing $n$ as the task complexity grows.
@@
A striking pattern emerges in the QRC performance for the SYK4 model [Fig. 3(c)]. When $\Delta t_{\mathrm{in}}$ is very small, the time evolution of the quantum reservoir does not yield an effective feature map, resulting in poor performance on both the STM and NARMA tasks. However, as $\Delta t_{\mathrm{in}}$ increases toward the Thouless time $t_{\mathrm{Th}}$, the QRC performance improves dramatically and reaches its maximum just before $t_{\mathrm{Th}}$. Beyond this point, both metrics level off at the baseline of the Haar QRC model for each order $n$, in accordance with the onset of universal RMT descriptions typical of quantum chaotic systems. In other words, the Thouless time effectively marks the temporal boundary where fully quantum chaotic signatures begin to dominate the QRC performance. The observed peak, situated just prior to $t_{\mathrm{Th}}$, thus explicitly underscores the performance enhancement at the “edge of many-body quantum chaos” in the temporal domain.
@@
In contrast, the QRC with the integrable SYK2 model behaves quite differently [Fig. 3(d)]. As $\Delta t_{\mathrm{in}}$ increases, the NARMA performance improves only modestly from its initial low values and quickly saturates. No pronounced performance peak emerges therein. This behavior starkly contrasts with that of the SYK4 model, indicating that the performance enhancements observed in Fig. 3(c) are fundamentally attributable to the latter’s intrinsic many-body quantum chaotic nature. Notably, despite its poor NARMA performance, the STM metric for the SYK2 model remains close to $1$ for sufficiently large values of $\Delta t_{\mathrm{in}}$, surpassing both the SYK4 model and the Haar QRC reference. This exemplifies the well-established trade-off in RC between memory capacity and nonlinear transformation capability [89, 90, 91]: QRC with the integrable SYK2 model retains strong memory but lacks nonlinear richness required for complex tasks. These model-dependent behaviors motivate our exploration of interpolations between the SYK2 and SYK4 models.
@@
Parametric edge of many-body quantum chaos. —We here demonstrate the QRC performance across the chaotic-integrable transition by mixing the SYK4 and SYK2 Hamiltonians. To ensure a fair comparison as the coupling strength is varied, each Hamiltonian is rescaled by its spectral norm: $\mathcal{H}\rightarrow\mathcal{H}/||\mathcal{H}||$ [92]. (Note that this correspondingly changes the time-scale relative to Fig. 3.) Figures 4(a) and 4(b) display the STM and NARMA performances versus the ratio $\kappa_{2}/J_{4}$ for long ($\Delta t_{\mathrm{in}}=50$) and short ($\Delta t_{\mathrm{in}}=1$) input intervals, respectively. The vertical dashed lines serve as a guide of the eye for the onset of WD and Poisson statistics, while the horizontal dotted lines represent the performance of the Haar QRC model for reference. We note that these input intervals are chosen to probe two distinct time-scales: the longer input interval $\Delta t_{\mathrm{in}}=50$ evaluates QRC performance within the universal RMT regime, while the shorter interval $\Delta t_{\mathrm{in}}=1$ characterizes QRC behavior in the nonuniversal, short-time window outside the scope of RMT (see the Supplemental Material [86]).

### v3

@@ Abstract
Reservoir computing (RC) is a machine learning paradigm that harnesses dynamical systems as computational resources. In its quantum extension—quantum reservoir computing (QRC)—these principles are applied to quantum systems, whose rich dynamics broadens the landscape of information processing. In classical RC, optimal performance is typically achieved at the “edge of chaos,” the boundary between order and chaos. Here, we identify its quantum many-body counterpart using the QRC implemented on the celebrated Sachdev-Ye-Kitaev model. Our analysis reveals substantial performance enhancements near two distinct characteristic “edges”: a temporal boundary defined by the Thouless time, beyond which system dynamics is described by random matrix theory, and a parametric boundary governing the transition from integrable to chaotic regimes. These findings establish the “edge of many-body quantum chaos” as a design guideline for QRC.
@@
More recently, the power of quantum has also integrated with reservoir computing (RC), a machine learning framework renowned for its efficacy in time-series analysis [12, 13, 14, 15, 16]. True to its name, the most distinctive feature of RC is the use of an input-driven dynamical system, known as a “reservoir”, to nonlinearly map input data onto a feature space. Notably, training in RC is limited to the post-processing stage, and the reservoir itself remains fixed, which significantly reduces the training cost. Extending this idea into the quantum domain, Fujii and Nakajima introduced quantum reservoir computing (QRC) [17]. QRC harnesses natural quantum dynamics as a reservoir, generating rich, nonlinear input-to-output mappings in an exponentially large Hilbert space without encountering the trainability bottlenecks prevalent in VQAs. This capability has sparked significant theoretical progress [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], as well as promising experimental demonstrations [39, 40, 41, 42, 43, 44, 45, 46, 47].
@@
Given the optimization-free nature of RC, researchers naturally ask: What characterizes an effective reservoir system? While this question remains largely unanswered for QRC, several key criteria have been identified for classical RC. The most renowned among them is the concept of the “edge of chaos” [48, 49, 50, 51, 52]. This posits that the boundary between chaotic and nonchaotic dynamical regimes provides a performance sweet spot for RC systems, as empirically verified across diverse platforms [53, 54, 55]. By analogy, exploring a quantum version of the edge of chaos offers a promising route to addressing the fundamental question posed above.
@@
In classical systems, chaos is naturally framed via phase space trajectories [56]. In quantum mechanics, however, the Heisenberg uncertainty principle precludes such a direct picture, and one must instead rely on more subtle diagnostics. Early investigations of quantum chaos therefore focused on single-particle systems whose classical limits are chaotic, leading to the conjecture (with some analytical supports) that their spectral statistics are governed by random matrix theory (RMT) [57, 58, 59, 60, 61]. In recent years, attention has shifted to many-body settings, spurred in large part by the exactly solvable Sachdev-Ye-Kitaev (SYK) model [62, 63, 64, 65]. Various diagnostics, including level statistics [66, 67, 68, 69], out-of-time-order correlators [70, 71, 72], and spectral form factors [73, 74, 75, 76] have enriched our understanding of chaos in interacting quantum systems. Experimental realizations of the SYK physics have also been proposed on several platforms [77, 78, 79, 80, 81]. Notably, the SYK model exhibits maximally quantum chaotic behavior, saturating the so-called bound on chaos [82, 72, 83] and has stood as a canonical model for the investigation of many-body quantum chaos. These insights motivate the exploration of QRC built upon the SYK model, aiming to address the open question of reservoir effectiveness from a quantum chaotic standpoint.
@@
In this Letter, we present a unified, RMT-based analysis that identifies the “edge of many-body quantum chaos” as the optimal guide for operating QRC. In particular, we examine the QRC performance near two distinct chaotic boundaries of the SYK model. The first is a temporal boundary defined by the Thouless time, beyond which the system’s dynamics conforms to RMT. The second is a parametric boundary marking the transition from quantum chaotic to integrable (nonchaotic) behavior, controlled by the strength of an additional noninteracting term. Using representative machine learning tasks that demand both memory retention and nonlinear transformation, we demonstrate that QRC performance is maximized near the onset of quantum chaotic regime in both temporal and parametric domains. These observations establish the “edge of many-body quantum chaos” as an important guideline for the design of QRC systems.
@@
where $N$ denotes the system size, and $c^{\dagger}_{i}$ ($c_{i}$) is the spinless fermionic creation (annihilation) operator at site $i$. In Eq. (1), the first term describes the SYK4 interaction: the all-to-all couplings $J_{ijkl}$ are drawn from a zero-mean complex Gaussian distribution of variance $\overline{|J_{ijkl}|^{2}}=J_{4}^{2}/N^{3}$ ($\overline{\bullet}$ denotes the ensemble average). The second term is the noninteracting SYK2 contribution, with coupling constants $\kappa_{ij}$ sampled from a zero-mean complex Gaussian distribution of variance $\overline{|\kappa_{ij}|^{2}}=\kappa_{2}^{2}/(2N)$. The couplings satisfy the symmetry relations $\kappa_{ij}=\kappa_{ji}^{*}$, $J_{ijkl}=J_{klij}^{*}$, and $J_{ijkl}=-J_{jikl}=-J_{ijlk}$. Both coupling terms conserve the total fermion number $N_{p}\equiv\sum_{i}\langle c_{i}^{\dagger}c_{i}\rangle$. Unless otherwise specified, we set $N=8$, $J_{4}=1$, and $N_{p}=N/2$ (note that the QRC input protocol introduced below does not conserve $N_{p}$).
@@
To diagnose the chaotic nature of quantum many-body systems, we first examine the level statistics [66, 67, 68, 69]. Let the ordered eigenvalues of $\mathcal{H}$ be $E_{n}$; the spacings between consecutive energy levels are given by $s_{n}=E_{n+1}-E_{n}$. According to RMT, the spacing distribution in integrable systems follows Poisson statistics, whereas in quantum chaotic systems, it is well described by Wigner-Dyson (WD) distribution. A convenient metric to characterize these distributions is the level spacing ratio, defined as $r_{n}=\min(s_{n}/s_{n+1},s_{n+1}/s_{n})$. Its mean value $\langle r\rangle$ assumes $\approx 0.3862$ for Poisson distribution and $\approx 0.5996$ for WD statistics [66, 67]. Figure 2(a) presents the distribution of the level spacing ratio $P(r)$ for the SYK4 model $(J_{4},\kappa_{2})=(1,0)$] and the SYK2 model $(J_{4},\kappa_{2})=(0,1)$], in excellent agreement with the WD and Poisson predictions, respectively. Notably, when both terms are present, the system can be tuned between these extremes: as the ratio of the typical coupling strengths $\kappa_{2}/J_{4}$ increases, Fig. 2(b) displays a gradual shift of $\langle r\rangle$ from the WD to the Poisson value. This result strongly suggests a transition from quantum-chaotic to integrable behavior [84, 85, 86, 87, 88]; for finite-size effects, see the Supplemental Material [89].
@@
where $\mathrm{Tr}_{1}\{\bullet\}$ denotes the partial trace over site $1$; see the Supplemental Material for details on the observed dynamics [89]. During the evolution, the expectation value of the single-site particle number is measured at discrete times $t_{v}=v\Delta t_{\mathrm{in}}/V$, where $v=1,\dots,V$ and $V$ is the number of virtual computational nodes (here, $V=10$); for the effect of statistical noise, see the Supplemental Material [89]. At the $k$-th step, the outcomes $\mathrm{Tr}\{\rho^{(k)}(t_{v})c_{i}^{\dagger}c_{i}\}$ are assembled into a one-dimensional state vector $\bm{\mathrm{x}}^{(k)}$; including an additional constant term, $\bm{\mathrm{x}}^{(k)}$ has $NV+1$ components in total. The output layer then applies a linear transformation $y^{(k)}=W_{\mathrm{out}}\bm{\mathrm{x}}^{(k)}$, where the output weight $W_{\mathrm{out}}$ is trained by minimizing the mean squared error between the computed output $y^{(k)}$ and the target output $\bar{y}^{(k)}$. In our simulations, a total of $10000$ inputs are processed. Starting from a random mixed state, the first $4000$ inputs are discarded to wash out the effects of initial conditions (see the Supplemental Material for the convergence property [89]). The subsequent $3000$ inputs are employed for training $W_{\mathrm{out}}$, and the final $3000$ are reserved for evaluation. Performance results are averaged over $500$ independent random realizations.
Benchmark tasks in reservoir computing are typically characterized by their requirements on temporal memory and nonlinear processing: the former specifies how far back in the input history the reservoir must retain information, while the latter is often described by the maximum polynomial order and the richness of cross-term interactions. Accordingly, we evaluate the reservoir performance on two complementary tasks: the short-term memory (STM) task, which probes linear memory capacity, and the nonlinear auto-regressive moving average (NARMA) task, which evaluates combined memory and nonlinear processing capabilities [14, 17]. Since real-world applications hinge on both properties, we place greater emphasis on the NARMA performance; for additional tasks, see the Supplemental Material [89].
@@
In the STM task, the target output at delay $d$ is defined as the input value $d$ steps earlier in a random input sequence $\{u^{(k)}\}$, i.e., $\bar{y}^{(k)}_{d}=u^{(k-d)}$. During the testing phase, we collect the outputs $\bm{\mathrm{y}}=\{y^{(k)}\}$ and targets $\bar{\bm{\mathrm{y}}}_{d}=\{\bar{y}^{(k)}_{d}\}$, and compute the determination coefficient $R^{2}_{d}\equiv\mathrm{cov}^{2}(\bm{\mathrm{y}},\bar{\bm{\mathrm{y}}}_{d})/[\sigma^{2}(\bm{\mathrm{y}})\sigma^{2}(\bar{\bm{\mathrm{y}}}_{d})]$, where $\mathrm{cov}$ and $\sigma^{2}$ denote the covariance and the variance, respectively. $R^{2}_{d}$ is bounded between 0 and $1$; values closer to $1$ indicate better memory performance.
In the NARMA task, the target sequence at a given order $n$ is defined by the recursive relation: $\bar{y}^{(k+1)}=\alpha\bar{y}^{(k)}+\beta\bar{y}^{(k)}\sum_{j=0}^{n-1}\bar{y}^{(k-j)}+\gamma u^{(k-n+1)}u^{(k)}+\delta$, where ($\alpha$, $\beta$, $\gamma$, $\delta$) are set to ($0.3$, $0.05$, $1.5$, 0.10.1[17, 20]. The input stream is sampled from the range $[0,0.2]$ to prevent divergence in $\{\bar{y}^{(k)}\}$, which is then rescaled to $[0,1]$ when encoded to the quantum reservoir [17]. Performance is evaluated in terms of the deviation between the output and the target, which is quantified by the normalized mean-squared error (NMSE) $\|\bm{\mathrm{y}}-\bar{\bm{\mathrm{y}}}\|^{2}/\|\bar{\bm{\mathrm{y}}}\|^{2}$ with $\|\bullet\|$ being the Euclidean norm. A lower NMSE indicates superior performance in accurately reproducing the target nonlinear dynamics.
@@
Temporal edge of many-body quantum chaos. —We begin with an examination of the QRC performance of the SYK4 and SYK2 models, in which clear fingerprints of many-body quantum chaos manifest themselves in the temporal domain. As a time-domain probe of the energy-spectrum correlations characteristic of quantum chaos, we analyze the spectral form factor (SFF) $K(t)=\overline{\sum_{m,n}e^{i(E_{m}-E_{n})t}}/\mathcal{N}^{2}$, where $\mathcal{N}$ denotes the Hilbert space dimension [73, 74, 75, 76]. In many-body chaotic systems, the SFF generally exhibits a “ramp-plateau” profile governed by two time-scales: the Thouless time $t_{\mathrm{Th}}$, after which the evolution follows RMT predictions, leading to a universal “ramp” reflecting level repulsion, and the plateau time $t_{\mathrm{p}}$, beyond which $K(t)$ saturates at a plateau due to the discreteness of the spectrum [90, 91, 92]. Figures 3(a) and 3(b) illustrate the SFF for the SYK4 and SYK2 limits, respectively. In the SYK4 case, the SFF exhibits the pronounced ramp beginning around $t_{\mathrm{Th}}\sim 4.5$ and the subsequent plateau beyond $t_{\mathrm{p}}\sim 80$, providing a clear signature of many-body quantum chaos. While $K(t)$ in the SYK2 model also features a dip and a plateau, they occur on a much shorter timescale and with a different system-size scaling of $t_{\mathrm{p}}$ compared to the chaotic system (see the Supplemental Material for details [89]).
@@
Figure 3: (a), (b) The SFF $K(t)$ for the SYK4 and SYK2 models, with each curve averaged over $20000$ realizations. (c), (d) The QRC performances as functions of $\Delta t_{\mathrm{in}}$ for the SYK4 and SYK2 models. The upper panels show the memory performance $R^{2}_{d=n-1}$ (higher values indicate better retention), while the lower panels display the NMSE on the NARMA tasks (lower values indicate better accuracy). Markers represent the orders $n=2,3,5$, and 77 (plotted sparsely for brevity). The horizontal dotted lines indicate the reference performance of the Haar QRC model, with poorer performance corresponding to higher $n$. Performance results are averaged over $500$ realizations, with shaded regions denoting the standard deviation. Vertical dashed lines in (a) and (c) represent the Thouless time $t_{\mathrm{Th}}$.
@@
Figures 3(c) and 3(d) plot the QRC performance of the SYK4 and SYK2 models as functions of the input interval $\Delta t_{\mathrm{in}}$, which defines the intrinsic time-scale governing the QRC dynamics. The upper panels display the STM performance $R^{2}_{d=n-1}$ at delay d=n1d=n-1, and the lower panels show the NMSE on the NARMA-$n$ tasks of orders $n=2$, 33, 55, and $7$. The chosen STM delay corresponds to the oldest input term explicitly appearing in the definition of the NARMA-$n$ dynamics; additional past inputs are nevertheless recursively encoded within the target sequence $\{\bar{y}^{k}\}$ (see the Supplemental Material for the full delay dependence of $R^{2}_{d}$ [89]). To establish a baseline for comparison, we also analyze the Haar QRC model, in which the evolution operator $e^{-i\mathcal{H}t}$ is replaced by a unitary $U$ drawn from the Haar measure; no intrinsic time scale exists therein [93, 94]. Its performance, shown as horizontal dashed lines, deteriorates monotonically with increasing $n$ as the task complexity grows.
@@
A striking pattern emerges in the QRC performance for the SYK4 model [Fig. 3(c)]. When $\Delta t_{\mathrm{in}}$ is very small, the time evolution operator of the quantum reservoir is nearly the identity, failing to enrich the input-to-output map and resulting in poor performance on both the STM and NARMA tasks. However, as $\Delta t_{\mathrm{in}}$ increases toward the Thouless time $t_{\mathrm{Th}}$, the QRC performance improves dramatically and reaches its maximum just before $t_{\mathrm{Th}}$. Beyond this point, both metrics level off at the baseline of the Haar QRC model for each order $n$, in accordance with the onset of universal RMT descriptions typical of quantum chaotic systems. In other words, the Thouless time effectively marks the temporal boundary where fully quantum chaotic signatures begin to dominate the QRC performance. The observed dip of the NMSE, situated just prior to $t_{\mathrm{Th}}$, thus explicitly underscores the performance enhancement at the “edge of many-body quantum chaos” in the temporal domain.
@@
In contrast, the QRC with the integrable SYK2 model behaves quite differently [Fig. 3(d)]. As $\Delta t_{\mathrm{in}}$ increases, the NARMA performance improves only modestly from its initial low values and quickly saturates, showing no pronounced peak. This behavior starkly contrasts with that of the SYK4 model, indicating that the performance enhancements observed in Fig. 3(c) are fundamentally attributable to the latter’s intrinsic many-body quantum chaotic nature. Notably, despite its poor NARMA performance, the SYK2 model attains higher STM performance than the SYK4 model. These model-dependent behaviors motivate our exploration of interpolations between the SYK2 and SYK4 models.
@@
Parametric edge of many-body quantum chaos. —We here demonstrate the QRC performance across the quantum chaotic-integrable transition by mixing the SYK4 and SYK2 Hamiltonians. To ensure a fair comparison as the coupling strength is varied, each Hamiltonian is rescaled by its spectral norm: $\mathcal{H}\rightarrow\mathcal{H}/$$\mathcal{H}||$ [95]. (Note that this correspondingly changes the time-scale relative to Fig. 3.) Figures 4(a) and 4(b) display the STM and NARMA performances versus the ratio $\kappa_{2}/J_{4}$ for long ($\Delta t_{\mathrm{in}}=50$) and short ($\Delta t_{\mathrm{in}}=1$) input intervals, respectively. The vertical dashed lines serve as a guide of the eye for the onset of WD and Poisson statistics, while the horizontal dotted lines represent the performance of the Haar QRC model for reference. We note that these input intervals are chosen to probe two distinct time-scales: the longer input interval $\Delta t_{\mathrm{in}}=50$ evaluates QRC performance within the universal RMT regime, while the shorter interval $\Delta t_{\mathrm{in}}=1$ characterizes QRC behavior in the nonuniversal, short-time window outside the scope of RMT.